In a geometric sequence whose first three terms are 9, -6, 4, ..., which term is 1024/6561?
Answer by Boreal(15235)

Consecutive terms have a ratio of -2/3, and the number of terms has to be even, because it is positive.
9(-2/3)^x=1024/6561=2^10/3^8. But we started with 9, or 3^2, so this is 10 terms
9*(-2/3)^10=1024/6561
10th term.